OpenStudy (anonymous):

the polynomial with real coefficient whose zeros are -4, 1-i, 1+ is ... a) x^3+6x^2+10x+8 b)x^3-2x^2-6x-8 c)x^3+2x^2-6x+8 d)x^3-6x^2+10x-8

OpenStudy (jdoe0001):

\(\large { x=\{4, 1-i, 1+ i\}\quad \\ \quad \\ \begin{cases} x=4\to x-4=0\to {\color{brown}{ (x-4)}}=0\\ x=1-i\to x-1+i=0\to {\color{brown}{ (x-1+i)}}=0\\ x=1+i\to x-1-i=0\to {\color{brown}{ (x-1-i)}}=0 \end{cases} \\ \quad \\ {\color{brown}{ (x-4)}}{\color{brown}{ (x-1+i)}}{\color{brown}{ (x-1-i)}}=0 \\ \quad \\ {\color{brown}{ (x-4)}}{\color{brown}{ (x-1+i)}}{\color{brown}{ (x-1-i)}}=\textit{original polynomial} }\)